# Stringlet Excitation Model of the Boson Peak

### v2

@@ V Outlook
$\frac{g_{s}(\omega_{BP})}{g_{Debye}^{T}(\omega_{BP})}\equiv v_{T}^{d}\frac{g_{s}(\omega_{BP})}{\omega_{BP}^{d-1}}\propto p(0)\lambda^{d+1},$ (28)
where $g_{Debye}^{T}(\omega_{BP})$ is the Debye contribution arising from the transverse acoustic modes $\omega^{d-1}/v_{T}^{d}$, up to some neglected constant numerical factors. It is interesting to notice that, in this approximation, the intensity is given uniquely in terms of the parameters appearing in the stringlet size distribution $p(l)$. Also, it is immediate to verify that the intensity of the BP as in Eq.(28) grows upon heating since the average length $\lambda$ increases with $T$.
Interestingly, the stringlet dynamics shares many similarities with the behavior of defects and anharmonic modes in crystals Granato (1992); Takeno and Sievers (1988) (See Yang, Wang, and Zaccone (2022) for a direct proof of the stringlet anharmonicity in relation to the BP in glasses.) and the role of weakly-dispersing soft optical modes, which has been identified in several instances as the origin for the BP anomaly in crystalline materials with no structural disorder Ackerman et al. (1981); Etrillard et al. (1996); Reményi et al. (2015); Cano and Levanyuk (2004); Moratalla et al. (2019); Takabatake et al. (2014); Schliesser and Woodfield (2015); Krivchikov et al. (2022); Baggioli and Zaccone (2019b). A similarity of this kind is also reflected by the fact that the string model of Lund was inspired by Granato and Lücke string model of crystal plasticity Granato and Lücke (1956). A unifying picture based on defect dynamics would be highly desirable and these collective excitations are a candidate for explaining the roton-like excitations observed previously in both liquids and glasses Grest, Nagel, and Rahman (1982); Suck et al. (1983); de Schepper et al. (1984), as reviewed in Astrath et al. (2006). Dispersion curves of this kind have also been suggested to be related to the BP anomaly in glasses Kovalenko and Krasny (1990); Krasny, Kovalenko, and Krawczyk (1998) (see also Zaccone and Baggioli (2022)), and might likewise be related to the stringlets.
Finally, our work revives the long-standing and vigorously debated question of whether the structures responsible for the BP are structural defects or dynamical excitations deriving from the anharmonicity of intermolecular interactions. The evidence would appear to favor the latter interpretation, explaining why the boson peak should also be expected in crystalline materials at elevated temperatures where anharmonicity in intermolecular interactions becomes prevalent as in cooled liquids. Recent simulations Zhang et al. (2019) have indeed confirmed that the boson peak also arises in crystalline materials at elevated temperatures, but below the melting temperature, where anharmonic intermolecular interactions start to become prevalent.
@@ Acknowledgements.
We would like to thank A. Zaccone, J. Zhang and M. A. Ramos for many discussions about the BP and the vibrational dynamics of amorphous solids. We are grateful to Y. Wang. and W. Xu for useful comments on a preliminary version of this draft. We thank P. Jacobson and C. Soles for helpful comments. We would like to thank H. Zhang for the picture in the right panel of Fig.1 and several fruitful discussions. CJ and MB acknowledge the support of the Shanghai Municipal Science and Technology Major Project (Grant No.2019SHZDZX01). MB acknowledges the sponsorship from the Yangyang Development Fund.
@@ Appendix A Analogy between dislocations in crystals and stringlets
Dislocations are extended “defect” structures that are prevalent in the modeling of the plasticity of crystalline materials. The model of “open” string-like defects utilized in the main text originates from a highly successful model of crystal plasticity introduced by Granato and Lücke Granato and Lücke (1956) to describe the “string-like” segments of the dislocations between pinning points. The development of this classical dislocation model of deformed crystalline materials is reviewed in Khonik (2021). In a more controversial work, Granato advocated that open string-like excitations in both crystalline materials and cooled liquids should correspond to extended defects involving “generalized” interstitial defects Granato (1992). Granato and his coworkers then advocated the radical idea that liquids should be viewed as crystalline materials populated by such defects. One of the many predictions of Granato’s model Granato (1992) strongly anticipates the results of the present paper. In particular, he argued that his interstitial defects should adopt a string like geometrical form and he “re-purposed” his former dislocation segment theory to model the boson peak of glass-forming liquids. This led him to a prediction of the boson peak in the low temperature limit of the same form as Eq.(10), although his admittedly rough derivation differs in the predicted prefactor. Moreover, his model made a rough estimate of the boson peak frequency. Clearly, our more refined model has much in common with the Granato model of the boson peak and the correspondence becomes even greater if an identity is granted between the strings and his hypothetical interstitial defects. This is a possibility that we plan to study in the future.
@@
While there is substantial experimental support for the ideas just described, it is admittedly unclear how one should interpret such defects in liquids (and amorphous systems in general) given there is no crystalline reference state that would allow for a discussion of well-defined structural defects, interstitial or otherwise. Nonetheless, both simulations and neutron scattering measurements have indicated that there are indeed “excitations” in the form of strings in the ps timescale of glass-forming liquids. Simulations show that such excitations also arise in model crystalline materials, although these structures cannot reasonably be identified as static dislocation loops or extended defects of composed of clustered interstitial defects Santos et al. (2018).
@@
It is our view that these string-like structures, that are notably observed in simulations of both crystalline and glass-forming materials, arise from the inherently anharmonic interactions in these materials, a phenomenon that is well-known in the Fermi-Ulam-Pasta-Tsingou model (e.g., Campbell (2004); Flach, Ivanchenko, and Kanakov (2006)). In particular, exact and numerical solutions of the dynamics and thermodynamics of such non-linear lattice models reveals the presence of intrinsically localized modes (sometimes termed “breathers”) Peyrard (1998); Peyrard and Farago (2000), deriving purely from the anharmonicity of the interparticle interactions.
@@
Importantly, the number of particles exhibiting this type of collective motion has been shown in the closely related $\phi^{4}$ lattice model to exhibit an exponential size distribution Flach and Mutschke (1994), as for stringlets. Similar distributions of the number of particles in this type on non-linear expectation have been noted in the analysis of the Peyrard-Bishop model of duplex DNA Peyrard and Bishop (1989), another closely related anharmonic 1-dimensional lattice model.
@@
Our theoretical framework adopted to describe the density of states distribution of the stringlets is inspired mathematically by the Granato-Lücke dislocation segment theory, noted above, and by a thermodynamic theory of string formation at equilibrium introduced previously to describe the $T$ dependence and length distribution of strings Betancourt P., Starr, and Douglas (2018). Very similar statistical mechanical models have been introduced to describe the formation of worm-like micelles, the equilibrium polymerization of S, P and other elements, etc.
@@
Despite the success of our modeling effort aimed at fundamentally understanding the boson peak, we must admit that the fundamental cause of this type of “excitation” is still not entirely clear. It is notable that string-like excitations are prevalent in many field theories of phase transitions and other thermodynamic transitions Copeland et al. (1991), such as the XY model Kohring, Shrock, and Wills (1986) and the superfluid to fluid transition in He4 whose critical properties are well-described by this model, vortex melting in type II superconductors Nguyen and Sudbø (1998), and cosmological models Zurek (1985); Antunes, Bettencourt, and Hindmarsh (1998); Antunes and Bettencourt (1998). We strongly believe that the formulation of such a gauge field theory of condensed matter along this line is required to address the fundamental origin of these dynamic excitations.

### v3

@@ V Outlook
$\frac{g_{s}(\omega_{BP})}{g_{Debye}^{T}(\omega_{BP})}\equiv v_{T}^{d}\frac{g_{s}(\omega_{BP})}{\omega_{BP}^{d-1}}\propto\mathcal{N}_{s}\lambda^{d+1},$ (30)
where $g_{Debye}^{T}(\omega_{BP})$ is the Debye contribution arising from the transverse acoustic modes $\omega^{d-1}/v_{T}^{d}$, up to some neglected constant numerical factors. It is interesting to notice that, in this approximation, the intensity is given uniquely in terms of the parameters appearing in the stringlet size distribution $p(l)$. Also, it is immediate to verify that the intensity of the BP as in Eq.(30) grows upon heating since $\langle l\rangle=\lambda+1$ increases with $T$.
Interestingly, the stringlet dynamics shares many similarities with the behavior of defects and anharmonic modes in crystals [71; 112] (See Ref.[113] for a direct proof of the stringlet anharmonicity in relation to the BP in glasses.) and the role of weakly-dispersing soft optical modes, which has been identified in several instances as the origin for the BP anomaly in crystalline materials with no structural disorder [114; 94; 93; 64; 6; 5; 115; 116; 65]. A similarity of this kind is also reflected by the fact that the string model of Lund was inspired by Granato and Lücke string model of crystal plasticity [56]. A unifying picture based on defect dynamics would be highly desirable and these collective excitations are a candidate for explaining the roton-like excitations observed previously in both liquids and glasses [117; 118; 119], as reviewed in [120]. Dispersion curves of this kind have also been suggested to be related to the BP anomaly in glasses [121; 122] (see also [123]), and might likewise be related to the stringlets.
Finally, our work revives the long-standing and vigorously debated question of whether the structures responsible for the BP are structural defects or dynamical excitations deriving from the anharmonicity of intermolecular interactions. The evidence would appear to favor the latter interpretation, explaining why the boson peak should also be expected in crystalline materials at elevated temperatures where anharmonicity in intermolecular interactions becomes prevalent as in cooled liquids. Recent simulations [96] have indeed confirmed that the boson peak also arises in crystalline materials at elevated temperatures, but below the melting temperature, where anharmonic intermolecular interactions start to become prevalent.
@@ Acknowledgements.
We would like to thank A. Zaccone, J. Zhang and M. A. Ramos for many discussions about the BP and the vibrational dynamics of amorphous solids. We are grateful to Y. Wang. and W. Xu for useful comments on a preliminary version of this draft. We thank P. Jacobson and C. Soles for helpful comments. We would like to thank H. Zhang for the unpublished image in the right panel of Fig.1 and several fruitful discussions. CJ and MB acknowledge the support of the Shanghai Municipal Science and Technology Major Project (Grant No.2019SHZDZX01). MB acknowledges the sponsorship from the Yangyang Development Fund.
@@ Appendix A Analogy between dislocations in crystals and stringlets
Dislocations are extended “defect” structures that are prevalent in the modeling of the plasticity of crystalline materials. The model of “open” string-like defects utilized in the main text originates from a highly successful model of crystal plasticity introduced by Granato and Lücke [56] to describe the “string-like” segments of the dislocations between pinning points. The development of this classical dislocation model of deformed crystalline materials is reviewed in [124]. In a more controversial work, Granato advocated that open string-like excitations in both crystalline materials and cooled liquids should correspond to extended defects involving “generalized” interstitial defects [71]. Granato and his coworkers then advocated the radical idea that liquids should be viewed as crystalline materials populated by such defects. One of the many predictions of Granato’s model [71] strongly anticipates the results of the present paper. In particular, he argued that his interstitial defects should adopt a string like geometrical form and he “re-purposed” his former dislocation segment theory to model the boson peak of glass-forming liquids. This led him to a prediction of the boson peak in the low temperature limit of the same form as Eq.(10), although his admittedly rough derivation differs in the predicted prefactor. Moreover, his model made a rough estimate of the boson peak frequency. Clearly, our more refined model has much in common with the Granato model of the boson peak and the correspondence becomes even greater if an identity is granted between the strings and his hypothetical interstitial defects. This is a possibility that we plan to study in the future.
@@
While there is substantial experimental support for the ideas just described, it is currently unclear how one should interpret such defects in liquids (and amorphous systems in general) given there is no crystalline reference state that would allow for a discussion of well-defined structural defects, interstitial or otherwise. Nonetheless, both simulations and neutron scattering measurements have indicated that there are indeed “excitations” in the form of strings in the ps timescale of glass-forming liquids. Simulations show that such excitations also arise in model crystalline materials, although these structures cannot reasonably be identified as static dislocation loops or extended defects of composed of clustered interstitial defects [125].
@@
It is our view that these string-like structures, that are notably observed in simulations of both crystalline and glass-forming materials, arise from the inherently anharmonic interactions in these materials, a phenomenon that is well-known in the Fermi-Ulam-Pasta-Tsingou model (e.g., [126; 127]). In particular, exact and numerical solutions of the dynamics and thermodynamics of such non-linear lattice models reveals the presence of intrinsically localized modes (sometimes termed “breathers”) [128; 129], deriving purely from the anharmonicity of the interparticle interactions.
@@
Importantly, the number of particles exhibiting this type of collective motion has been shown in the closely related $\phi^{4}$ lattice model to exhibit an exponential size distribution [130], as for stringlets. Similar distributions of the number of particles in this type on non-linear expectation have been noted in the analysis of the Peyrard-Bishop model of duplex DNA [131], another closely related anharmonic 1-dimensional lattice model.
@@
Our theoretical framework adopted to describe the density of states distribution of the stringlets is inspired mathematically by the Granato-Lücke dislocation segment theory, noted above, and by a thermodynamic theory of string formation at equilibrium introduced previously to describe the $T$ dependence and length distribution of strings [37]. Very similar statistical mechanical models have been introduced to describe the formation of worm-like micelles, the equilibrium polymerization of S, P and other elements, etc.
@@
Despite the success of our modeling effort aimed at fundamentally understanding the boson peak, we must admit that the fundamental cause of this type of “excitation” is still not sufficiently clear. It is notable that string-like excitations are prevalent in many field theories of phase transitions and other thermodynamic transitions [132], such as the XY model [133] and the superfluid to fluid transition in He4 whose critical properties are well-described by this model, vortex melting in type II superconductors [134], and cosmological models [135; 136; 137]. We strongly believe that the formulation of such a gauge field theory of condensed matter along this line is required to address the fundamental origin of these dynamic excitations.